Based on the below data what will be the value of mse?

Regression Statistics
Multiple R
R Square
Adjusted R Square
Standard Error
Observations 8
ANOVA
df SS MS F
Regression 1 23 23.0 11.5
Residual 6 12 2.0
Total 7
Coefficients Standard Error t Stat P-value
Intercept 20 31.274666 3.984284 0.007248
Advertising (thousands of $) 41 6.19330674 1.610802 0.158349
Answer #1
Regression Statistics
Multiple R
R Square
Adjusted R Square
Standard Error
Observations 8
ANOVA
degree of freedom sum of square mean square F
Regression 1 =(k) 23 =(RSS) 23.0 =(MSR) 11.5 =(MSR/MSE)
Residual 6 (n-k-1) 12 (SSE) 2.0 = MSE = (SSE/n-k-1)
Total 7 (n-1)
Coefficients Standard Error t Stat P-value
Intercept 20 31.274666 3.984284 0.007248
Advertising (thousands of $) 41 6.19330674 1.610802

0.158349

Here, MSE is already calculated.

It is SSE/(n-k-1) which is 2.